OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

isnt it multiple choice

OpenStudy (texaschic101):

2(x- 1) >= - 12 2x - 2 >= -12 2x >= -12 + 2 2x >= -10 x >= -5 or -5 <= x closed circle on -5 (because there is an equal sign) and shading to the right (greater then shaded to the right). These are the correct ways.